OpenStudy (anonymous):

a painter leans a 20 ft ladder against a building. the base of the ladder is 12 ft from the building. from the nearest foot, how high from the building does the ladder reach?

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

c=20ft

OpenStudy (anonymous):

12^2+b^2=20^2

OpenStudy (lizz123):

\[a^{2}+b^{2}=c^{2}\]

OpenStudy (anonymous):

sqrt(144-400)=b